OpenStudy (anonymous):

i need help finding the zeros of this equation help please? x3 – x2 + 5x – 5 = 0

OpenStudy (anonymous):

x^3 - x^2 + 5x = 0

OpenStudy (anonymous):

3x2-5x-2=0 Find the factors such that the product of the factors is the trinomial 3x2-5x-2. This can be done by trial and error and checked using the FOIL method of simplifying polynomials. (3x+1)(x-2)=0 Set each of the factors of the left-hand side of the equation equal to 0. 3x+1=0 Set each of the factors of the left-hand side of the equation equal to 0. x-2=0 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. 3x=-1 Divide each term in the equation by 3. 3x/3=-1/3 Simplify the left-hand side of the equation by canceling the common terms. x=-1/3 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. x=2 The complete solution is the set of the individual solutions. x=-1/3,2 hope this helps best of luck and dont forget to choose me for the best answer because this took me so long